OpenStudy (anonymous):

Provide the complete solution to: 3x - 2 ≡ 7 (mod 11) 5 ≡ 4x - 1 (mod 9??)

myininaya (myininaya):

\[3x-2 \equiv 7 (\mod 11) => 3x-2-7=11k \text{ for some integer k}\] \[3x-9=11k\] \[3x=11k+9\] k=0 => x=3 k=1 => x=not integer k=2 => x=not integer k=3 => x=14 k=4 => x=not integer k=5 => x=not integer k=6 => x=25 k=7 => x=not integer k=8 => x=not integer x=9 => x=36 so we have x=3,14,25,36,... x=11i+3 for i=0,1,2,3,4,....

myininaya (myininaya):

above just say where i is integer you know so we can include all integers not just the positives \[5 \equiv 4x-1 (\mod 9) =>5-(4x-1) = 9k \text{ for some integer k}\] \[-4x+6 =9k\] \[4x=6-9k\] k=-2=>x=6 k=-1=>x=not integer k=0=> x=not integer k=1=> x=not integer k=2=> x=-3 k=3=> x=not integer k=4=> x=not integer k=5=> x=not integer k=6=> x=-12 k=7=> x=not integer k=8=> x=not integer k=9=> x=not integer k=10=>x=-21 k=11=>not integer x=6,-3,-12,-21,... x=42-9(4+i) where i is integer